OpenStudy (anonymous):

Use a Double- or Half-Angle Formula to solve the equation in the interval. Help please! tan θ/2 − sin θ = 0

OpenStudy (campbell_st):

well if you use half angle then its basically \[\tan(\frac{\theta}{2}) = t\] and \[\sin(\theta) = \frac{2t}{1 + t^2}\] it comes from the trig identity  and the fact that \[\sin(\theta) = 2\sin(\frac{\theta}{2})\cos(\frac{\theta}{2})\] so making the subsitutions you get \[t - \frac{2t}{\sqrt{1 + t^2}} = 0\] so you just need to solve for t. hope it helps